Find : a)the bank discount, and b) the proceeds if Nick settles his debt of $3700 at the end of 3 months at an annual discount rate of 13%.
Answer #1

Part (a) Bank Discount = Principal x Discount rate x time = $ 3,700 x 13% x 3 / 12 = $ 120.25

Part (b) Proceeds = Principal - bank discount = 3,700 - 120.25 = $ 3,579.75